OpenStudy (angelwings996):

f(x) = 3x - 7 and g(x) = -2x - 6. Find (f o g)(4) Please help!?

OpenStudy (hba):

First plug in g(x)=-2x-6 in f(x)=3x-7 in place of x :)

OpenStudy (angelwings996):

So would it be f(x) = 3(-2x - 6) - 7 ?

OpenStudy (hba):

Yes f(g(x))=3(-2x-6)-7 Now put 4 in place of x and get (fog)(4)=.. or f(g(4))=..

OpenStudy (angelwings996):

Okay, so would it be f(g(4)) = -49 ?

OpenStudy (hba):

Yeah right :)

OpenStudy (angelwings996):

Okay Thank you! could you help me with one more please ?

OpenStudy (hba):

Sure :)

OpenStudy (angelwings996):

Okay, Rewrite \[\sqrt[3]{27x - 81} - 5\] to make it easy to graph using a translation. Describe the graph

OpenStudy (angelwings996):

I know that you would put 1/3 as an exponent but I don't know where

OpenStudy (hba):

Well see this. \[\huge\ \sqrt[3]{a}=(a)^{1/3}\] Another example, \[\huge\ \sqrt[3]{x-2}=(x-2)^{1/3}\]

OpenStudy (angelwings996):

Okay so it would be (27x - 81)^1/3 - 5, right ?

OpenStudy (hba):

Yes we can work more on it if you want :)

OpenStudy (angelwings996):

Yes please! Would this be the equation to make it easier to graph or is there more ?

OpenStudy (hba):

Yes lets continue then take the 27 common. \[\huge\ [27(x-1)]^{1/3}-5 \] \[\huge\ [3^3 (x-1)]^{1/3}-5\]

OpenStudy (hba):

Now you can cancel. \[\huge\ (x^3)^{(1/3)}=x\]

OpenStudy (angelwings996):

Okay, is this what the equation is or is this just stating what x equals ?

OpenStudy (hba):

I am showing you how you can cancel.

OpenStudy (hba):

The above statement was an example.

OpenStudy (angelwings996):

Oh okay

OpenStudy (hba):

\[\huge\ [3^3 (x-1)]^{1/3}-5\] Can you further simplify this ?

OpenStudy (angelwings996):

Wait, I am not sure how you did the (x^3) ^1/3 = x

OpenStudy (angelwings996):

Would it be (9x - 1)^1/3 - 5 ? @hba

OpenStudy (hba):

No.

OpenStudy (angelwings996):

I'm confused :/

OpenStudy (hba):

\[\huge\ (3^3)^{1/3} ((x-1))^{1/3}-5\]Now can you do it ?

OpenStudy (angelwings996):

Lets see... (3^3)^1/3 would become just 3 correct?

OpenStudy (hba):

Yeah :)

OpenStudy (angelwings996):

Okay so it would be 3(x - 1) - 5 ? or would I combine 3 and x - 1?

OpenStudy (hba):

No. 3 (x-1)^{1/3}-5

OpenStudy (angelwings996):

Oh, I forgot the 1/3 sorry

OpenStudy (angelwings996):

Then what would I do?

OpenStudy (hba):

Have some ice-cream :) Can i get some cookies ? I am the cookie monster.

OpenStudy (hba):

C is for cookie and cookie is for me. P.S. I am out of cookies :'(

OpenStudy (angelwings996):

Wait, is that the answer ?

OpenStudy (hba):

Yeah LoL

OpenStudy (angelwings996):

Oh haha, then I would just graph it and describe what it looks like right ?

OpenStudy (angelwings996):

I've got some cookies lol (:

OpenStudy (hba):

cookies LoL :P Yeah you can graph it :) Now it is easy i guess :)

OpenStudy (angelwings996):

Okay lol, Thank you for your help!

OpenStudy (hba):

Welcome Welcome :)